# A Quantitative Evaluation of the Expressivity of BMI, Pose and Gender in Body Embeddings for Recognition and Identification

### v1

@@ V-C Advantages of Expressivity for Person ReID
    Enables Attribute Comparison Across Features: In the context of person ReID, comparing the significance of various attributes within a network’s feature set is critical for understanding how identity information is organized. Expressivity facilitates this by normalizing the error rates of attributes to a common scale, as illustrated in Figures 5 and 6. This standardization enables direct comparison of attribute contributions, unlike conventional metrics influenced by varying scales.
@@
    Supports Both Discrete and Continuous Attributes: This work focuses on body attributes for ReID, expressivity is versatile and can be applied to both discrete (e.g., gender) and continuous (e.g., pitch angle) attributes. For instance, the expressivity of a concept such as gender can be computed if a binary attribute vector $A$ (indicating 1 for male and 0 for female or vice versa) is available. Unlike TCAV [24], which is well-suited for discrete attributes, expressivity extends to continuous concepts like pose angle or BMI, where identifying negative examples (images where the concept is absent) is challenging. This capability is particularly valuable for person ReID, where continuous attributes often play a critical role.
  3. 3.
@@
VI Conclusion
We propose a method to quantify the information a ViT-based person ReID network learns about various attributes without being explicitly trained on them, by analyzing their expressivity on learnt features. This enables us to identify attributes most relevant to identity recognition across hierarchical layers and training epochs. Several important findings emerge from our investigation: (1) BMI consistently shows the highest expressivity, especially in deeper layers (e.g., layer 12) and later training stages (e.g., epoch 11), making it the most critical attribute for identity recognition even without explicit labels. (2) Attributes like yaw and pitch are expressive in mid-layers (e.g., layers 4 and 6) but lose influence in deeper layers. (3) Temporally, BMI expressivity increases throughout training, while yaw and pitch decline sharply, with yaw showing the steepest drop. Gender, notably, has minimal correlation with learned features. These findings highlight BMI as the most significant attribute, followed by yaw and pitch and gender for the person ReID task. However, since expressivity approximates MI, it is influenced by entropy and attribute label distribution, potentially affecting cross-attribute comparisons which is an inherent limitation of all MI-based approaches.
VII ACKNOWLEDGMENTS
@@
VIII Ethical Impact Statement
This research utilizes the BRIAR dataset, which was collected by the Test and Evaluation team under strict consent guidelines and with approval from the appropriate ethical review board. All subjects involved provided informed consent for their participation, including the use of their images in research publications and figures. Our study aims to analyze how demographic and physiological covariates such as body mass index and gender may introduce biases in body-based recognition systems. By systematically examining these potential biases, we seek to contribute to the broader discourse on fairness, accountability, and transparency in biometric technologies. Understanding and mitigating bias in recognition systems is critical for ensuring equitable outcomes across diverse populations. This research aligns with the outlined ethical principles by advocating for responsible AI development and fostering inclusivity in biometric system evaluation.

### v3

@@ 5.3 Advantages of Expressivity for Person ReID
    Supports Both Discrete and Continuous Attributes: Expressivity is versatile and applicable to both discrete (e.g., gender) and continuous (e.g., pitch angle) attributes. For example, gender expressivity can be computed using a binary attribute vector $A$. Unlike TCAV [24], which is tailored to discrete attributes, expressivity naturally extends to continuous concepts such as pose or BMI where defining clear negative examples is difficult. This makes it especially valuable in ReID, where continuous attributes are often crucial.
@@
6 Conclusion
We propose a method to quantify the information a ViT-based person ReID network learns about various attributes without being explicitly trained on them, by analyzing their expressivity on learnt features. This enables us to identify attributes most relevant to identity recognition across hierarchical layers and training epochs. Several important findings emerge from our investigation: (1) BMI consistently shows the highest expressivity, especially in deeper layers (e.g., layer 12) and later training stages (e.g., epoch 11), making it the most critical attribute for identity recognition even without explicit labels. (2) Attributes like yaw and pitch are expressive in mid-layers (e.g., layers 4 and 6) but lose influence in deeper layers. (3) Temporally, BMI expressivity increases throughout training, while yaw and pitch decline sharply, with yaw showing the steepest drop. Gender is entangled in the features moderately and has minimal variation with the evolution of learned features. These findings highlight BMI as the most significant attribute, followed by yaw gender and pitch for the person ReID task. However, since expressivity approximates MI, it is influenced by entropy and attribute label distribution, potentially affecting cross-attribute comparisons which is an inherent limitation of all MI-based approaches.
7 Acknowledgements